OpenStudy (anonymous):

What is the solution to the equation 9^(x + 2) = 27?

OpenStudy (anonymous):

both sides are powers of 3

OpenStudy (anonymous):

\[9^{x+2}=(3^2)^{x+2}=3^{2x+4}\] and \[27=3^3\] which means you have \[3^{2x+4}=3^3\] and so \[2x+4=3\] now you can solve for \(x\)